# CleanVul: Automatic Function-Level Vulnerability Detection in Code Commits Using LLM Heuristics

### v1

@@ 2. Motivating Example
Consider the following real-world example [^2] https://github.com/thingsboard/thingsboard-edge/commit/1a3ee8512d58625940b25e46bc6488a3539fdc5e from the ThingsBoard project, which illustrates the complexity of identifying vulnerability fixes within commits. The commit message is presented below, targeting XSS vulnerabilities. This commit introduces security measures such as NoXss validation to prevent XSS attacks, yet it also encompasses a variety of modifications unrelated to direct vulnerability mitigation. These include enhancing the codebase by removing unnecessary imports and updating license documentation, as well as improving test coverage to confirm the efficacy of the new validations.
@@ 5.2. Model Choices
Our experiments encompass various state-of-the-art LLMs. We have selected three widely-used encoder-only models, namely RoBERTa (Liu et al., 2019), CodeBERT (Feng et al., 2020), and GraphCodeBERT (Guo et al., 2020), along with three decoder-only models, which consist of two smaller models, GPT-2 (Radford et al., 2019) and CodeGPT (Zan et al., 2022), and one larger decoder-only model, CodeLlama (Roziere et al., 2023). These models have demonstrated exceptional performance in code-related tasks, making them suitable candidates for our study (Xu et al., 2022).
@@ 6.1. RQ1: What is the correctness improvement of CleanVul compared to the uncleaned dataset?
To evaluate Correctness of CleanVul, we randomly select a sample of 487 code changes from the collected GitHub VFC dataset and manually analyze them to create a testing dataset. We carefully examine each code change and mark them as either vulnerability-fixing changes or other code changes, establishing a ground truth for our evaluation. This evaluation method aligns with previous work (Ding et al., 2024). As mentioned in Section 3, VulSifter’s output score ranges from 0 to 4, with 0 indicating no relation to vulnerability fixes and 4 signifying a strong focus on fixing vulnerabilities. We construct four datasets by setting different threshold values. A very clean dataset includes only items scored as 4, representing the highest confidence in vulnerability-fixing changes. In contrast, a more extensive dataset retains items scored 1 or higher, encompassing potential vulnerability-fixing changes. The results are presented in Table 3.
@@
Regarding generalization to PrimeVul, training CodeGPT on CleanVul on Java and testing on PrimeVul resulted in an accuracy of 55.13%, which is very close to the best performance achieved by training and testing GraphCodeBERT on PrimeVul (55.26%). Conversely, when training on PrimeVul and testing on CleanVul on Java, the best performance was 53.91% for CodeBERT, which is considerably lower than the 74.78% accuracy achieved when training and testing on CleanVul (Java only). Additionally, the highest accuracy obtained when training on SVEN and testing on PrimeVul was 53.73% using GPT-2 or RoBERTa, which is lower than the results from training on CleanVul on Java and testing on PrimeVul. This evidence highlights the high quality of our dataset, suggesting that CleanVul (Java only) may contain more comprehensive or diverse examples of vulnerability-fixing changes than PrimeVul and SVEN, and that the knowledge can be generalized to PrimeVul but not vice versa.
@@ 10. Conclusion and Future Work
We created CleanVul, a new high-quality vulnerability dataset, by analyzing 5,352,105 commits from 127,063 GitHub repositories. Using VulSifter, we filtered out noise from vulnerability-fixing commits (VFCs) in this corpus and applied different confidence thresholds to balance dataset size and quality. With a threshold of 3, CleanVul contains 11,632 vulnerable function pairs achieving 90.6% Correctness, while a stricter threshold of 4 yields 8,337 function pairs with 97.3% Correctness in our test sample. These results are comparable to established datasets such as SVEN (94.0% Correctness) and PrimeVul (86.0% Correctness), while overcoming their limitations. Unlike SVEN, which relies on manual analysis and contains only 803 samples, our approach is scalable. Additionally, while PrimeVul requires NVD entry links, our VulSifter can analyze all VFCs, including those without such links. Our evaluation of various LLMs fine-tuned on CleanVul revealed its superior generalization capabilities across different datasets. Notably, models fine-tuned on CleanVul significantly outperformed PrimeVul-trained models when tested on SVEN, achieving 64.87% accuracy compared to 55.75%. When evaluated on PrimeVul, our models demonstrated better performance than those trained on SVEN (58.09% vs 55.20%). Most remarkably, models trained on CleanVul and tested on PrimeVul achieved higher accuracy than models trained and tested on PrimeVul itself (58.09% vs 57.19%). These results underscore CleanVul’s effectiveness and robustness for model fine-tuning, particularly in its ability to capture generalizable vulnerability patterns that transfer well across different contexts.

### v2

@@ 2. Motivating Example
Consider the following real-world example [^2] https://github.com/thingsboard/thingsboard-edge/commit/1a3ee8512d58625940b25e46bc6488a3539fdc5e from the ThingsBoard project, which illustrates the complexity of identifying vulnerability fixes within commits. The commit message is presented below, targeting XSS vulnerabilities. This commit exemplifies what researchers call tangled commits (Kirinuki et al., 2014) - commits that address multiple concerns or include changes serving different purposes simultaneously. While the primary purpose is introducing security measures such as NoXss validation to prevent XSS attacks, the commit also encompasses unrelated modifications including enhancing the codebase by removing unnecessary imports and updating license documentation, as well as improving test coverage to confirm the efficacy of the new validations.
@@ 5.2. Model Choices
Our experiments encompass various state-of-the-art LLMs. We have selected three widely-used encoder-only models, namely RoBERTa (Liu et al., 2019), CodeBERT (Feng et al., 2020), and GraphCodeBERT (Guo et al., 2020), along with three decoder-only models, which consist of two smaller models, GPT-2 (Radford et al., 2019) and CodeGPT (Lu et al., 2021), and one larger decoder-only model, CodeLlama (Roziere et al., 2023). These models have demonstrated exceptional performance in code-related tasks, making them suitable candidates for our study (Xu et al., 2022).
@@ 6.1. RQ1: What is the correctness improvement of CleanVul compared to the uncleaned dataset?
To evaluate Correctness of CleanVul, we randomly select a sample of 487 function-level code changes from the collected GitHub VFC dataset and manually analyze them to create a testing dataset. This sample size provides a confidence level of 95% with a margin of error of ±4.4% for the full dataset, allowing us to make statistically significant observations. We carefully examine each code change and mark them as either vulnerability-fixing changes or other code changes, establishing a ground truth for our evaluation. This evaluation method aligns with previous work (Ding et al., 2024). As mentioned in Section 3, VulSifter’s output score ranges from 0 to 4, with 0 indicating no relation to vulnerability fixes and 4 signifying a strong focus on fixing vulnerabilities. We construct four datasets by setting different threshold values. A very clean dataset includes only items scored as 4, representing the highest confidence in vulnerability-fixing changes. In contrast, a more extensive dataset retains items scored 1 or higher, encompassing potential vulnerability-fixing changes. The results are presented in Table 3.
@@
Regarding generalization to PrimeVul, training CodeGPT on CleanVul on Java and testing on PrimeVul resulted in an accuracy of 55.13%, which is very close to the best performance achieved by training and testing GraphCodeBERT on PrimeVul (55.26%). Conversely, when training on PrimeVul and testing on CleanVul on Java, the best performance was 53.91% for CodeBERT, which is considerably lower than the 74.78% accuracy achieved when training and testing on CleanVul (Java only). Additionally, the highest accuracy obtained when training on SVEN and testing on PrimeVul was 53.73% using GPT-2 or RoBERTa, which is lower than the results from training on CleanVul on Java and testing on PrimeVul. This evidence highlights the high quality of our dataset, suggesting that CleanVul (Java only) may contain more comprehensive or diverse examples of vulnerability-fixing changes than PrimeVul and SVEN, and that the knowledge from training on CleanVul can be generalized to PrimeVul but not vice versa.
@@ 10. Conclusion and Future Work
We created CleanVul, a new high-quality vulnerability dataset, by analyzing 5,352,105 commits from 127,063 GitHub repositories. Using VulSifter, we filtered out noise from vulnerability-fixing commits (VFCs) in this corpus. VulSifter has a configurable threshold that can be set based on data cleanliness requirements, allowing us to balance dataset size and quality. With a threshold of 3, CleanVul contains 11,632 vulnerable function pairs achieving 90.6% Correctness, while a stricter threshold of 4 yields 8,337 function pairs with 97.3% Correctness in our test sample. These results are comparable to established datasets such as SVEN (94.0% Correctness) and PrimeVul (86.0% Correctness), while overcoming their limitations. Unlike SVEN, which relies on manual analysis and contains only 803 samples, our approach is scalable. Additionally, while PrimeVul requires NVD entry links, our VulSifter can analyze all VFCs, including those without such links. Our evaluation of various LLMs fine-tuned on CleanVul revealed its superior generalization capabilities across different datasets. Notably, models fine-tuned on CleanVul significantly outperformed PrimeVul-trained models when tested on SVEN, achieving 64.87% accuracy compared to 55.75%. When evaluated on PrimeVul, our models demonstrated better performance than those trained on SVEN (58.09% vs 55.20%). Most remarkably, models trained on CleanVul and tested on PrimeVul achieved higher accuracy than models trained and tested on PrimeVul itself (58.09% vs 57.19%). These results underscore CleanVul’s effectiveness and robustness for model fine-tuning, particularly in its ability to capture generalizable vulnerability patterns that transfer well across different contexts.